### v4

# PaperIIPaper $\defcitealiasPaperIIIPaper II$I

Abstract
DIPLODOCUS (Distribution-In-PLateaux methODOlogy for the CompUtation of transport equationS) is a framework being developed for the mesoscopic modelling of astrophysical systems via the transport of particle distribution functions through the seven dimensions of phase space, including continuous forces and discrete interactions between particles. This first paper in a series provides an overview of the analytical framework behind the model, consisting of an integral formulation of the relativistic transport equations (Boltzmann equations) and a discretisation procedure for the particle distribution function (Distribution-In-Plateaux). The latter allows for the evaluation of anisotropic interactions and generates a conservative numerical scheme for a distribution function’s transport through phase space.
1 Introduction
Astronomical phenomena must usually be inferred by distant observation alone. Blazars are a class of objects whose observed emissions are thought to originate from magnetised, relativistic jets of material formed close to the black holes at the centre of Active Galactic Nuclei (AGN) and directed almost exactly along our line of sight. The spectra of such objects are truly multi-wavelength, spanning from radio to ultra-high-energy gamma rays, typically with a double-peaked shape (Urry and Padovani, 1995; Padovani et al., 2017; Blandford et al., 2019; Hovatta and Lindfors, 2019).
The processes that dominate this emission are generally thought to be synchrotron radiation and inverse Compton scattering from either leptonic or hadronic populations, or a combination of both, within the jets. The observed brightness of such processes is heavily dependent on the properties of the in situ particle populations; it is therefore of critical importance that such processes and populations are well modelled.
In deriving macroscopic (hydrodynamic) models such as (General Relativistic Magneto-)HydroDynamics (GRM)HD, particle populations are assumed to be highly collisional and in local thermal equilibrium. Therefore, knowledge of their distribution in momentum space may be neglected in favour of a small set of defining parameters, e.g. temperature, density and bulk flow velocity. This reduction of dimensionality permits large spatial domains to be modelled, leading to high resolution simulations of the global jetted AGN structure (McKinney and Gammie, 2004; McKinney, 2006; Chatterjee et al., 2019; Whitehead and Matthews, 2023) at the cost of limited information on the underlying micro-scale interactions between particle populations, hence poor direct simulation of emissions.
By contrast, microscopic (kinetic) models maintain information about the underlying distribution of particles in momentum space, at the cost of the increased dimensionality. To make this computationally tractable, other assumptions, e.g. isotropy in momentum, are often made to reduce the dimensionality as much as possible.
One frequently used microscopic approach, in the field of plasma physics, is that of Particle-In-Cell (PIC). PIC replaces particle distributions with a finite number of discrete particles that attempt to simulate the entire distribution. This approach works well when distributions are near thermal or when small scale plasma instabilities and dynamics require resolution, such as for particle acceleration in magnetised shock fronts and reconnection regions (Cerutti et al., 2013; Sironi et al., 2015; Mehlhaff et al., 2024; Comisso et al., 2024). However, due to computational limits, PIC can only simulate a finite number of particles. This limits sampling of both the distribution and interaction processes and constrains the spatial extent of any simulation. Therefore, PIC cannot typically be applied on a macroscopic scale.
An alternative microscopic approach is to assume symmetries in the particle distribution in space and momentum space. For example, an emitting region of an astrophysical source can be simplified to a spherical or slab-like zone with an internally isotropic distribution of particles. Early astrophysical use of these methods was in modelling cosmic ray transport (Lindquist, 1966; Webb, 1985, 1989). Here, in addition to spherical symmetry in momentum space, specifically with reference to a co-moving frame, particle distributions were taken to be close to Maxwellian. The assumption of isotropy in particular led to extensive work (for example Blumenthal and Gould, 1970; Weaver, 1976; Aharonian and Atoyan, 1981; Svensson, 1982; Baring, 1987; Dermer, 1984; Brinkmann, 1984; Dermer, 1985, 1986; Coppi and Blandford, 1990; Sarkar et al., 2019) describing a wide range of interactions between particles and fields, with only small ventures into the world of anisotropy for specific (Stepney and Guilbert, 1983; Moskalenko and Strong, 2000; Kelner et al., 2014; Lai and Ng, 2023) and general (Everett and Cotter, 2024) interactions. These works form the foundation of modern single/multi-zone models for jetted astrophysical sources (Mastichiadis and Kirk, 1995; Mastichiadis et al., 2005; Dimitrakoudis et al., 2012; Potter and Cotter, 2012, 2013a, 2013b, 2013c; Böttcher et al., 2013; Cerruti et al., 2015; Gao et al., 2017; Potter, 2018; Fernández and van Eerten, 2021; Lucchini et al., 2022; Boula and Mastichiadis, 2022; Zacharias et al., 2022; Hahn et al., 2022; Klinger et al., 2024a; Stathopoulos et al., 2024; Cerruti et al., 2024), which maintain the tradition of momentum-space isotropy to limit computational requirements. These requirements are further reduced by considering additional symmetries in the spatial distribution of particles. In general, such models operate by considering independent, homogeneous zones of isotropic plasma, normally spherical or slab-like in shape, with prescribed internal properties such as bulk Lorentz factor, magnetic field strength and initial particle distributions. These zones may then be placed along the axis of the jet to generate its structure, though they typically do not interact with one another. The distributions of particles within each zone are then evolved according to their interactions and the parameters of the zone, with emission spectra generated by Doppler boosting escaping photons from each zone into an observer frame. These spectra can then be fit to observational data of astrophysical sources. These zoned model fits have provided great insight into the environment and dynamics of sources like AGN jets, allowing the study of questions such as how particles are accelerated to the initial distributions, and whether the jet material is dominated by leptonic or hadronic particles, questions remain a matter of debate (Weekes, 2003; Hoerbe et al., 2020; Kantzas et al., 2023).
The insight these zoned models have provided into blazars, and jetted astrophysical sources in general, such as X-Ray Binaries (XRB) and Gamma-Ray Bursts (GRBs), should not be understated (see, for example Markoff et al., 2001; Böttcher et al., 2013; Marscher, 2014; Potter and Cotter, 2015; Potter, 2018; Kantzas et al., 2021; Klinger et al., 2024b). However, zoned models are intrinsically limited by their constraining symmetries and it is yet to be seen if their results are robust upon relaxation of these assumptions.
Recent studies in the fields of pair-plasma winds (Aksenov et al., 2004, 2007, 2010) and cosmic-ray propagation (Krumholz et al., 2022, 2024) have demonstrated self-consistent evolution of axisymmetric particle distributions using microscopic approaches. Therefore, the additional dimensionality that comes with relaxing isotropy is within the capability of modern computing. Further, the study of the globular cluster Terzan 5 by Krumholz et al. (2024) provided a unique explanation for inconsistent gamma-ray emissions. Low-energy gamma rays are observed coincident with the location of the globular cluster, while a source of high-energy gamma rays is spatially displaced in the direction of Terzan 5’s magnetotail. This displacement was successfully modelled by the slow isotropisation of an accelerated population of cosmic rays propagating down the magnetotail. Initially, due to their narrow range of pitch angles, their emissions are beamed away from the line of sight and only once the distribution becomes sufficient isotropic do the emissions become visible. This explanation of the observed spatial separation was only made possible by considering the anisotropic evolution of the cosmic-ray population.
These findings raise many questions regarding the assumed momentum-space isotropy and spatial homogeneity of zoned models. If these assumptions were to be eliminated, would the simulated emissions continue to fit the observed data? Would this require a reinterpretation of the source, and could this provide deeper physical insights? These questions motivated the present work, which derives and details a framework designed to address them: DIPLODOCUS.
With the elimination of assumed symmetries, kinetic models boil down to a question of particle transport: how does a particle go from point $A$ to point $B$? (With the slight complication that point $A$ and point $B$ exist not in three dimensions of space but the seven dimensions of phase space that include space, time and momentum.)
This general approach to particle transport includes aspects of both microscopic and macroscopic models; small-scale interactions and large-scale structure. As such it lies between the two—in the mesoscopic regime. This regime, and the generality of the framework, permit its application to whole host of astrophysical phenomena, not just jetted sources.
[…]
2 Coordinate-Free Transport Equation
This section defines (with some rigour) a coordinate-free, covariant form of the transport equations for particle distribution functions through phase space, also more generally known as Boltzmann equations (Boltzmann, 1872). The use of a coordinate-free approach is mathematically compact, and allows the transport equations to be adapted to any physical (or even non-physical) systems.
In the context of astrophysical sources, particles encounter continuous conservative (Hamiltonian) forces, such as the electromagnetic Lorentz force, as well as non-conservative (non-Hamiltonian) forces, such as radiation reaction that generates synchrotron emission. Particles may also encounter discrete, discontinuous interactions, i.e. collisions, which affect their path through phase space.
There is a rich history of developing Boltzmann equations for the purpose of astrophysical modelling. However, most derivations focus solely on conservative cases, neglecting non-conservative elements. This work primarily builds on the approaches taken by Lindquist (1966); Ehlers (1971); Groot et al. (1980). Other works, some derived from the above sources and some independent, that have made impressions on the work presented here include Bichteler (1965); Hakim (1967a, b); Stewart (1971); Steeb (1979, 1980); Ellis et al. (1983a, b); Blandford and Eichler (1987); Cercignani (2002); Cardall and Mezzacappa (2003); Lifshits (2008); Shibata et al. (2014).
2.1 A to B: Phase Flow on a Manifold
As a first step, consider a single particle located at some point $A$ at some instance and at $B$ at a later instance. These points $A$ and $B$ exist in an abstract phase space, a differentiable manifold of dimension $n$, denoted by $\mathcal{M}$. Both points exist along a unique curve $\gamma(\lambda)$, where $\lambda$ is an affine parameter along that curve such that $\gamma(\lambda_{A})=A$ and $\gamma(\lambda_{B})=B$. The coordinates of any point along this curve have the form $\bm{y}=\left(y^{1},...,y^{n}\right)$ and there exists a unique vector $\bm{v}$ tangent to the curve at each point. The curve $\gamma(\lambda)$ is therefore defined by the equation of motion
$\frac{\mathrm{d}\bm{y}}{\mathrm{d}\lambda}=\bm{v},$ (1)
and is known as the particle’s worldline through phase space (see Fig. 1).
Figure 1: The worldline of a single particle through phase space ($\mathcal{M}$) between two points $A$ and $B$, defined by a curve $\gamma(\lambda)$, with $\lambda$ being some affine parameter of that curve, and tangent to the vector $\bm{v}$.
Now consider expanding this concept to every point $q\in\mathcal{M}$, i.e. at every point there exists a vector field $\bm{v}$ which defines a unique congruence of worldlines $\gamma_{q}(\lambda)$ that fill $\mathcal{M}$. These curves define how any particle will evolve if placed at any point $q\in\mathcal{M}$ (see Fig. 2).
The manifold $\mathcal{M}$ has a volume form (volume element) associated with it defined by
$\bm{\Omega}=\Omega(\bm{y})\bm{\mathrm{d}}y^{1}\wedge...\wedge\bm{\mathrm{d}}y^{n}=\Omega(\bm{y})\bm{\mathrm{d}}y^{1...n},$ (2)
where $\wedge$ is the wedge product and $\bm{\mathrm{d}}y^{1...n}$ is shorthand for $\bm{\mathrm{d}}y^{1}\wedge...\wedge\bm{\mathrm{d}}y^{n}$. The evolution of this volume form with the phase flow generated by $\bm{v}$ is given by a Lie derivative with respect to $\bm{v}$ (also known as the Liouville operator or vector)
$\mathcal{L}_{\bm{v}}\bm{\Omega}=\bm{\mathrm{d}}(\bm{v}\lrcorner\bm{\Omega})+\bm{v}\lrcorner(\bm{\mathrm{d}}\bm{\Omega}),$ (3)
[…]
$\mathcal{L}_{\bm{v}}\bm{\Omega}=\bm{\mathrm{d}}(\bm{v}\lrcorner\bm{\Omega})=(\text{div}_{\bm{\Omega}}\bm{v})\bm{\Omega}=\bm{\mathrm{d}}\bm{\omega},$ (4)
where $\text{div}_{\bm{\Omega}}$ is the $\bm{\Omega}$-divergence. Note that $\bm{v}\lrcorner\bm{\Omega}\equiv\bm{\omega}$ defines the hypersurface element for hypersurfaces in $\mathcal{M}$.
Figure 2: A congruence of worldlines $\gamma_{q}$ through phase space $\mathcal{M}$ defined by the vector field $\bm{v}$ at every point $q\in\mathcal{M}$.
2.2 The Distribution Function
For Hamiltonian (conservative) systems, $\bm{\mathrm{d}}\bm{\omega}=0$ and therefore $\mathcal{L}_{\bm{v}}\bm{\Omega}=0$, which is a statement of Liouville’s theorem (Liouville, 1838) that the phase space volume is invariant with respect to the phase flow. Systems of importance to astrophysics are not in general Hamiltonian (see Section 3.1) and as such the phase space volume is not, in general, invariant with respect to the phase flow i.e. $\mathcal{L}_{\bm{v}}\bm{\Omega}\neq 0$. Hence Liouville’s theorem does not apply and instead must be generalised.
Consider a new volume form $\bar{\bm{\Omega}}\equiv f(\bm{y})\bm{\Omega}$, where $f$ is some differentiable scalar field on $\mathcal{M}$, defined such that the new phase space volume element satisfies:
$\mathcal{L}_{\bm{v}}\bar{\bm{\Omega}}=\mathcal{L}_{\bm{v}}(f\bm{\Omega})=0.$ (5)
The condition on $f$ for such a statement to be true may be obtained by expanding the Lie derivative using the product rule,
$\mathcal{L}_{\bm{v}}(f\bm{\Omega})=\left(\bm{v}(f)+f\text{div}_{\bm{\Omega}}\bm{v}\right)\bm{\Omega}=0.$ (6)
For non-trivial $\bm{\Omega}$, Eq. (6) provides a differential transport equation for $f$ through phase space
$\bm{v}(f)+f\text{div}_{\bm{\Omega}}\bm{v}=0.$ (7)
Further, the function $f$ may be identified by considering Stokes’ theorem applied to an arbitrary region of phase space $Q\in\mathcal{M}$:
$\int_{Q}\mathcal{L}_{\bm{v}}(f\bm{\Omega})=\int_{Q}\bm{\mathrm{d}}(f\bm{\omega})=\int_{\partial Q}f\bm{\omega}=0,$ (8)
where $\partial Q$ is the bounding surface of $Q$. The measure $\int_{\partial Q}f\bm{\omega}$ can then freely be identified with the number of worldlines (particles) passing through the closed hypersurface $\partial Q$, and $f\bm{\omega}$ the flux of particles through $\partial Q$. Hence Eq. (8) simply states that the divergence in phase space of the particle flux is equal to zero, i.e. the number of worldlines (particles) entering the region $Q$ is equal to the number leaving (see Fig. 3). Therefore $f$ may be defined as the phase space density of worldlines (particles) on hypersurfaces in $\mathcal{M}$ or, more colloquially, the particle distribution function.
Figure 3: Introduction of the distribution function $f$, a scalar field on the manifold $\mathcal{M}$ along with an integration domain $Q$ with volume element $\bm{\Omega}$ and boundary $\partial Q$ with hypersurface element $\bm{\omega}$. The integral $\int_{\partial Q}f\bm{\omega}$ then measures the number of worldlines (particles) passing through the hypersurface $\partial Q$.
2.3 Adding Collisions
The differential and integral transport equations for the distribution function, Eqs. (7) and (8), assume an absence of discrete, discontinuous interactions (collisions) within $\mathcal{M}$ between particles. A collision at a point $q\in\mathcal{M}$ may be defined as a termination or starting of a worldline (more precisely the occupation of a worldline by a particle) due to a particle being destroyed/created or receiving a discrete transfer of momentum. For example, a particle $\mathfrak{a}$ [^3] Particle species are referred to by Gothic scripts in this paper. at a position $q_{\mathfrak{a}}$ in its phase space $\mathcal{M}_{\mathfrak{a}}$ may receive a discrete quantity of momentum due to scattering off a particle $\mathfrak{b}$ at position $q_{\mathfrak{b}}\in\mathcal{M}_{\mathfrak{b}}$. The scattering causes the termination of the worldlines of particles $\mathfrak{a}$ and $\mathfrak{b}$ at points $q_{\mathfrak{a}}$ and $q_{\mathfrak{b}}$ respectively as each particle’s momentum (a coordinate in phase space) has changed by a discrete amount. The worldlines for particles $\mathfrak{a}$ and $\mathfrak{b}$ are then re-started at some new positions $q^{\prime}_{\mathfrak{a}}$ and $q^{\prime}_{\mathfrak{b}}$ as shown in Fig. 4. Alternatively, particle $\mathfrak{a}$ may undergo some emissive or decay process, where again it will undergo some discrete change in momentum and therefore terminate and re-start its worldline, while also starting new worldlines for any secondary particles emitted. Interactions, scattering, collisions, emissive and decay process all have the same effect in phase space, that being the termination and starting of worldlines, hence the term collision will be used to refer collectively to these processes.
These collisions can then be represented by a new measure $\bm{C}(\bm{y})\propto\bm{\Omega}$ known as the collision integral, which “counts” the number of collisions within the volume $Q$ (see further in Section 3.3). This measure can therefore be simply added to the right hand side of Eqs. (7) and (8), to give the differential transport equations for $f$:
$\bm{v}(f)+f\text{div}_{\bm{\Omega}}\bm{v}=\bm{C},$ (9)
and its integral form:
$\int_{\partial Q}f\bm{\omega}=\int_{Q}\bm{C}.$ (10)
Figure 4: Collisions (discrete transfers of momentum) are described by the termination (crosses) and starting (bar) of particle worldlines within phase space. Here a particle $\mathfrak{a}$ at a position $q_{\mathfrak{a}}$ collides with a particle $\mathfrak{b}$ at a position $q_{\mathfrak{b}}$, terminating each worldline. The locations $q_{\mathfrak{a}}$ and $q_{\mathfrak{b}}$ need not be coincident as the particles may have different momenta (a coordinate in phase space). The collision may cause a discrete exchange of momentum, whereby the particles then re-appear at points $q^{\prime}_{\mathfrak{a}}$ and $q^{\prime}_{\mathfrak{b}}$, starting a new pair of worldlines. The integral $\int_{Q}\bm{C}$ then counts the number of collisions (worldlines that are started and/or terminated) within the volume $Q$.
3 Coordinate-Dependent Transport Equation
3.1 Phase Flow with Coordinates
Though Eqs. (9) and (10), are mathematically elegant in their coordinate-free form, some definition of coordinates (basis vectors) are needed to give them utility in modelling astrophysical sources.
For this purpose, two sets of basis vectors are used. The nature of spacetime is assumed to be a 4D Riemannian manifold $\mathcal{N}$, with signature (+++)(-+++), and a coordinate basis $\bm{e}_{\alpha}=$ (Greek scripts) with coordinates $x^{\alpha}=\{x^{0},...,x^{3}\}$ [^4] The choice of a precise set of coordinates, e.g. Cartesian, cylindrical, spherical, Schwarzschild, Boyer-Lindquist, etc. remains arbitrary. As such this model can be applied to a wide range of spacetimes.. For momentum space, which exists in the tangent space $T_{\mathcal{N}}$ to $\mathcal{N}$, a local-orthonormal basis (tetrad) $\bm{e}_{a}$ (Latin scripts), with components $p^{a}=\{p^{0},...,p^{3}\}$ is used. The worldline of a particle through the 8-dimensional phase space is then given by
$\frac{\mathrm{d}x^{\alpha}}{\mathrm{d}\lambda}=p^{a}e^{\alpha}_{a},\quad\frac{\mathrm{d}p^{a}}{\mathrm{d}\lambda}=-\Gamma^{a}_{~bc}p^{b}p^{c}+mF^{a},$ (11)
where the transformation between the two basis is $\bm{e}_{a}=e_{a}^{~\alpha}\bm{e}_{\alpha}$, $\Gamma^{a}_{~bc}$ are the connection coefficients in the local orthonormal basis (i.e. Ricci rotation coefficients) and $F^{a}$ are the components of some arbitrary four-force $\bm{F}$. For massive particles $m\neq 0$ the affine parameter $\lambda$ is related to the proper time $\tau$ by $m\lambda=\tau$. The manifold $\mathcal{M}$, corresponding to the 8-dimensional phase space, is given by the set
$\begin{split}\mathcal{M}=&\{(\bm{x},\bm{p}):\bm{x}\in\mathcal{N},\bm{p}\in\bm{T}_{\mathcal{N}},\\ &\quad\bm{p}^{2}\leq 0,\bm{p}\text{ future directed}\}.\end{split}$ (12)
The phase flow in $\mathcal{M}$ is generated by the Lie derivative with respect to $\bm{v}$, where, using Eq. (11)
$\bm{v}=p^{a}e_{a}^{~\alpha}\frac{\partial}{\partial x^{\alpha}}+\left(-\Gamma^{a}_{~bc}p^{b}p^{c}+mF^{a}\right)\frac{\partial}{\partial p^{a}}.$ (13)
The rest mass of a particle
$m^{2}=\bm{p}^{2}=g_{\alpha\beta}e^{\alpha}_{~a}e^{\beta}_{~b}p^{a}p^{b},$ (14)
is a scalar function on $\mathcal{M}$ which, on physicality grounds, should be a constant with respect to the phase flow i.e.
$\mathcal{L}_{\bm{v}}(-m^{2})=0.$ (15)
Via direct calculation [^5] This requires the additional relation between the connection coefficients in a orthonormal basis $\Gamma^{a}_{~bc}$ (Ricci rotation coefficients) and those in a coordinate basis $\Gamma^{\alpha}_{~\beta\gamma}$ (Christoffel symbols): $\Gamma^{a}_{~bc}=e^{a}_{~\alpha}e_{c}^{~\gamma}\left(\partial_{\gamma}e_{b}^{~\alpha}+\Gamma^{\alpha}_{~\beta\gamma}e_{b}^{~\beta}\right).$ (16) , this yields the following constraint on the force $\bm{F}$:
$\bm{F}\cdot\bm{p}=0,$ (17)
which is met for all forces expected to occur within an astrophysical setting (e.g. Lorentz and radiation reaction forces). With $m^{2}$ being constant with respect to the phase flow, particles exist only on hypersurfaces in $\mathcal{M}$ defined by $m^{2}=\text{const}$. This restricted manifold is now denoted $\mathcal{M}_{m}$ and is 7-dimensional with coordinates $x^{\alpha}=\{x^{0},...,x^{3}\}$ and $p^{i}=\{p^{1},...,p^{3}\}$, where second-half Latin alphabet letters refer to indices valued $i\in\{1,2,3\}$, where $p^{0}$, the “energy” of the particle, has been chosen to be the dependent variable defined by
$p^{0}=\sqrt{m^{2}+(p^{i})^{2}}.$ (18)
The phase flow on the restricted manifold $\mathcal{M}_{m}$ is then given by
$\bm{v}=p^{a}e_{a}^{~\alpha}\frac{\partial}{\partial x^{\alpha}}+\left(-\Gamma^{i}_{~bc}p^{b}p^{c}+mF^{i}\right)\frac{\partial}{\partial p^{i}}.$ (19)
What are the volume forms on $\mathcal{M}_{m}$? The spacetime volume form is given by the standard form [^6] The Hodge dual of the unit function $\bm{\chi}=\star 1$.
$\bm{\chi}=\frac{\chi_{\alpha\beta\gamma\delta}(\bm{x})}{4!}\bm{\mathrm{d}}x^{\alpha\beta\gamma\delta},$ (20)
where $\chi_{\alpha\beta\gamma\delta}$ are components of the completely antisymmetric Levi-Civita tensor, with $\chi_{1234}=\sqrt{-\det(g_{\alpha\beta})}=\sqrt{-g}$. The momentum space volume, restricted to $\mathcal{M}_{m}$, is likewise given by
$\bm{\pi}=2\Theta(p^{0})\delta(\bm{p}^{2}-m^{2})\sqrt{-\eta}\bm{\mathrm{d}}p^{0123}=\frac{\sqrt{-\eta}}{p^{0}}\bm{\mathrm{d}}p^{123},$ (21)
where $\eta_{ab}$ is the metric in momentum space. For an orthonormal basis $\sqrt{-\eta}=\sqrt{-\det(\eta_{ab})}=1$, hence
$\bm{\pi}=\frac{1}{p^{0}}\bm{\mathrm{d}}p^{123}=\frac{\pi_{ijk}}{3!}\bm{\mathrm{d}}p^{ijk}.$ (22)
The total volume form is then given by the wedge product of Eqs. (20) and (22):
$\bm{\Omega}=\bm{\chi}\wedge\bm{\pi}.$ (23)
From Eq. (23), the hypersurface element $\bm{\omega}$ orthogonal to the phase flow $\bm{v}$ (Eq. (19)) in $\mathcal{M}_{m}$ can also be evaluated:
$\begin{split}\bm{\omega}\equiv\bm{v}\lrcorner\bm{\Omega}=&p^{a}e_{a}^{~\alpha}\frac{\chi_{\alpha\beta\gamma\delta}(\bm{x})}{3!}\bm{\mathrm{d}}x^{\beta\gamma\delta}\wedge\bm{\pi}\\ &+\bm{\chi}\wedge\left(-\Gamma^{i}_{~bc}p^{b}p^{c}+mF^{i}\right)\frac{\pi_{ijk}}{2!}\bm{\mathrm{d}}p^{jk}.\end{split}$ (24)
As discussed in Section 2.2, for a system to obey Liouville’s theorem $\bm{\mathrm{d}}\bm{\omega}=0$ is required, by consideration of Eq. (24) this condition is equivalent to $\frac{\partial F^{a}}{\partial p^{a}}=0$. As an example where this condition is met, the Lorentz force on a particle of charge $q$ and mass $m$ is given by $F^{a}=\frac{q}{m}F^{ab}p_{b}$, where $F^{ab}$ is the electromagnetic field tensor, therefore $\frac{\partial F^{a}}{\partial p^{a}}=F^{a}_{~a}=0$ as $F^{ab}$ is antisymmetric. However, a counter example is the radiation reaction force [^7] The radiation reaction force is also known as the Abraham-Lorentz-Dirac force (Abraham, 1905; Lorentz, 1892; Dirac, 1938) (see also Gralla et al., 2009; Poisson et al., 2011, for more modern descriptions). experienced by a charged particle being accelerated by an electromagnetic field where $F^{a}\propto\left(\eta^{ab}+\frac{p^{a}p^{b}}{m^{2}c^{2}}\right)\left(\frac{q}{m}p^{d}p^{c}\partial_{c}F_{bd}+\frac{q^{2}}{m^{2}}qF_{bc}F^{cd}p_{d}\right)$, for such a force it can be shown that $\frac{\partial F^{a}}{\partial p^{a}}\neq 0$, and is therefore a non-conservative force, evident by the fact that it results in the charged particle losing energy over time.
As such non-conservative forces, e.g. radiation reaction, are a necessary inclusion of any self-consistent astrophysical model; hence $\mathbf{d}\bm{\omega}\neq 0$ in general and the correct forms for the transport equation for the particle distribution function are Eqs. (9) and (10).
3.2 Stationary Observers
It is important to define a set of observers with respect to whom the distribution function and its derived properties are measured. If the spacetime is stationary, it permits a set of Eulerian observers. Their worldlines are defined everywhere by the timelike unit vector $\bm{n}=n_{\alpha}\bm{e}^{\alpha}=n_{0}\bm{\mathrm{d}}t$, with $\bm{n}\cdot\bm{n}=n^{0}n_{0}=-1$. At each value of proper time $\tau$ along the worldline of an observer, there exists a unique, spacelike hypersurface $\Sigma_{\tau}$ orthogonal to the worldline. This hypersurface defines the set of events that are considered simultaneous with respect to that observer at that value of proper time.
The hypersurface element $\bm{\omega}$, given by Eq. (24), may be decomposed into hypersurface elements of simultaneous and non-simultaneous events as measured by a stationary observer:
$\begin{split}\bm{\omega}&=\left[(-p^{a}n_{a})n^{\alpha}\frac{\chi_{\alpha\beta\gamma\delta}(\bm{x})}{3!}\bm{\mathrm{d}}x^{\beta\gamma\delta}\wedge\bm{\pi}\right]\\ &\quad+\left[\left(p^{a}e^{~\alpha}_{a}+p^{a}n_{a}n^{\alpha}\right)\frac{\chi_{\alpha\beta\gamma\delta}(\bm{x})}{3!}\bm{\mathrm{d}}x^{\beta\gamma\delta}\wedge\bm{\pi}\right.\\ &\quad+\left.\bm{\chi}\wedge\left(-\Gamma^{i}_{~bc}p^{b}p^{c}+mF^{i}\right)\frac{\pi_{ijk}}{2!}\bm{\mathrm{d}}p^{jk}\right]\\ &=\bm{\omega}_{\Sigma}+\bm{\omega}_{\Lambda}.\end{split}$ (25)
The first square bracketed term in Eq. (25) is $\bm{\omega}_{\Sigma}$, referring to the element of $\Sigma$ hypersurfaces, while the second square bracket is the element of timelike hypersurfaces $\Lambda$ of non-simultaneous events, $\bm{\omega}_{\Lambda}$ (see Fig. 5).
Figure 5: Timelike $\Sigma$ and spacelike $\Lambda$ hypersurfaces in phase space $\mathcal{M}_{m}$ defined by a set of stationary observers with trajectory tangent to $\bm{n}$. The spacelike hypersurfaces $\Sigma$ define a set of simultaneous events as measured by the observer.
With a set of observers, the total number of particles in the system can be counted in a well defined way. Consider the insight from Section 2.2, that the $\int_{\partial Q}f\bm{\omega}$ is the number of worldlines passing through some hypersurface defined by the boundary $\partial Q$. Therefore, taking $\partial Q$ to be the open hypersurface $\Sigma$, $\int_{\Sigma}f\bm{\omega}=N$ the number of particles measured simultaneously passing through that surface, i.e. the total number of particles within the manifold at a given time $\tau$ according to that local observer. If $\Sigma=X\times P$ is decomposed into the product of a spatial submanifold $X$ and the momentum-space submanifold $P$ this integral can be expanded as:
$\begin{split}N&=\int_{\Sigma}f\bm{\omega}_{\Sigma}\\ &=\int_{X}-n_{a}n^{\alpha}\frac{\chi_{\alpha\beta\gamma\delta(\bm{x})}}{3!}\bm{\mathrm{d}}x^{\beta\gamma\delta}\int_{P}fp^{a}\bm{\pi}.\end{split}$ (26)
The latter part of Eq. (26) can be identified as the first moment of the distribution function i.e. locally measured 4-flow vector:
$N^{a}=\int_{P}fp^{a}\bm{\pi},$ (27)
from which the scalar number density of particles $n$ as measured by a stationary observer can be defined:
$n=-n_{a}N^{a}.$ (28)
Higher moments of the distribution function can similarly be defined, for example the locally measured stress-energy tensor is defined by
$T^{ab}=\int_{P}fp^{a}p^{b}\bm{\pi},$ (29)
through which the scalar energy density $e=n_{a}n_{b}T^{ab}$ and pressure $p=\frac{1}{3}T^{ab}\Delta_{ab}$ (where $\Delta_{ab}=\eta_{ab}+n_{a}n_{b}$ is the projection tensor) can be defined.
An alternative to the set of Eulerian observers is a set of observers co-moving with some bulk flow. However, for co-moving observers, their local tetrad $\bm{e}_{a}$ is a function of the bulk flow, and therefore variable, unlike a Eulerian observer whose tetrad is fixed. This results in the hypersurface elements Eq. (24) also becoming a function of the flow and therefore so too would the numerical fluxes (see Section 4.2) used to evaluate transport of the distribution function, thereby requiring unnecessary recalculation when the bulk flow changes.
3.3 The Collision Integral
The collision integral $\bm{C}(\bm{x},\bm{p})$ is a measure of the number of worldlines that are terminated and started within some volume of phase space. It is useful to relate this to the transition probability $w_{\mathfrak{if}}$ for an initial state $\mathfrak{i}$ of particles to transition to a final state $\mathfrak{f}$. Modifying the results of Berestetskii et al. (1982), the differential transition probability may be written:
$\bm{w}_{\mathfrak{if}}=\delta^{(4)}(\bm{p}_{\mathfrak{i}}-\bm{p}_{\mathfrak{f}})\left|T_{\mathfrak{if}}\right|^{2}\bm{\chi},$ (30)
where the Dirac delta function describes conservation of momentum during the transition and $T_{\mathfrak{if}}$ is the standard scattering $T$-matrix.
Given a distribution of particles $f(\bm{x},\bm{p})$, the number of particle worldlines crossing a $\Sigma$ hypersurface is given by Eq. (26). Assuming a well defined basis transformation $n_{a}=(-1,0,0,0)$ and Eq. (26) may be integrated over a test region in the spatial sub-manifold $X$, giving
$\int_{X}\bm{\omega}_{\Sigma}=p^{0}V\bm{\pi},$ (31)
where $V$ is some spatial volume. As such the expression $p^{0}V\bm{\pi}$ can be thought of as the momentum-space number density of particles with a unity distribution function, i.e. 1 particle per unit hypersurface element. The differential transition probability, Eq. (30), may then be expanded to include the incident and final particle’s momentum states:
$\bm{w}_{\mathfrak{if}}=\delta^{(4)}(\bm{p}_{\mathfrak{i}}-\bm{p}_{\mathfrak{f}})\frac{\left|T_{\mathfrak{if}}\right|^{2}}{\mathfrak{n}_{\mathfrak{i}}!\mathfrak{n}_{\mathfrak{f}}!}\bm{\chi}\prod_{\mathfrak{a}\in\mathfrak{i}}p^{0}_{\mathfrak{a}}V\bm{\pi}_{\mathfrak{a}}\prod_{\mathfrak{b}\in\mathfrak{f}}p^{0}_{\mathfrak{b}}V\bm{\pi}_{\mathfrak{b}},$ (32)
where $\mathfrak{n}_{\mathfrak{i}}$ and $\mathfrak{n}_{\mathfrak{f}}$ are the number of identical initial and identical final states respectively, which are included to avoid over counting states.
In the literature, the factors of $p^{0}V$ are often absorbed into the scattering matrix $T_{\mathfrak{if}}$ to generate the more familiar scattering matrix $M_{\mathfrak{if}}$ i.e.
$M_{\mathfrak{if}}=T_{\mathfrak{if}}\sqrt{\prod_{\mathfrak{a}\in\mathfrak{i}}p^{0}_{\mathfrak{a}}V\prod_{\mathfrak{b}\in\mathfrak{f}}p^{0}_{\mathfrak{b}}V},$ (33)
giving
$\bm{w}_{\mathfrak{if}}=\delta^{(4)}(\bm{p}_{\mathfrak{i}}-\bm{p}_{\mathfrak{f}})\frac{\left|M_{\mathfrak{if}}\right|^{2}}{\mathfrak{n}_{\mathfrak{i}}!\mathfrak{n}_{\mathfrak{f}}!}\bm{\chi}\prod_{\mathfrak{a}\in\mathfrak{i}}\bm{\pi}_{\mathfrak{a}}\prod_{\mathfrak{b}\in\mathfrak{f}}\bm{\pi}_{\mathfrak{b}}.$ (34)
The number of worldlines terminated or started by an collision $\mathfrak{i}\rightarrow\mathfrak{f}$ is then given by reintroducing non-unity distribution functions for the initial particles:
$\bm{N}_{\mathfrak{if}}=\delta^{(4)}(\bm{p}_{\mathfrak{i}}-\bm{p}_{\mathfrak{f}})\frac{\left|M_{\mathfrak{if}}\right|^{2}}{\mathfrak{n}_{\mathfrak{i}}!\mathfrak{n}_{\mathfrak{f}}!}\bm{\chi}\prod_{\mathfrak{a}\in\mathfrak{i}}f_{\mathfrak{a}}(\bm{x},\bm{p})\bm{\pi}_{\mathfrak{a}}\prod_{\mathfrak{b}\in\mathfrak{f}}\bm{\pi}_{\mathfrak{b}}.$ (35)
With the volume element for phase space given by Eq. (23), it can be recognised that the quantity $\bm{N}_{\mathfrak{if}}$ is proportional to $\bm{\Omega}$, and therefore related to the collision integral $\bm{C}(\bm{x},\bm{p})$. To understand this relation, consider a particle of type-$\mathfrak{a}$. The number of its worldlines terminated by collisions where it is a member of the incident state $\mathfrak{i}$ is given by $\sum_{\mathfrak{a}\in\mathfrak{i}}\bm{N}_{\mathfrak{if}}$, and similarly the number of its worldlines that are started is $\sum_{\mathfrak{a}\in\mathfrak{f}}\bm{N}_{\mathfrak{if}}$.
The nature of the collision integral for a single particle of type-$\mathfrak{a}$ is therefore
$\bm{C}(\bm{x},\bm{p}_{\mathfrak{a}})=\sum_{\mathfrak{a}\in\mathfrak{f}}\bm{N}_{\mathfrak{if}}-\sum_{\mathfrak{a}\in\mathfrak{i}}\bm{N}_{\mathfrak{if}}.$ (36)
Specific examples of the collision integral for binary (scattering) collisions i.e. $\mathfrak{ab}\rightarrow\mathfrak{cd}$ and emission (decay) processes i.e. $\mathfrak{a}\rightarrow\mathfrak{bc}$ are given in Appendix A.
Eq. (36) contains integrals over the momentum states of all the initial and final state particles; it is therefore of high dimensionality and further complicated by potentially non-linear dependence on the particle distribution functions. As such, it is generally impractical to use in this form unless some methodology for simplifying the integral is applied, or some assumptions are made for the symmetries of $f$ in momentum space—traditionally, isotropy.
4 Distribution-In-Plateaux
Figure 6: The continuous distribution function $f$ for each particle is discretised via Eq. (37) to form a range of plateaux over phase space (here taken to be two dimensional with coordinates $x$ and $y$) indexed by the values $f_{ij}$, hence the name Distribution-In-Plateaux.
The integrated form of the Boltzmann equations (Eq. (10)) is mathematically elegant but in its present form is difficult to evaluate. The difficulty arises from the integrals being evaluated over the entire volume of phase space. This can be simplified by splitting the integration domain $Q$ into a set of continuous sub-domains with common boundaries. Therefore the left hand side of Eq. (10) describes the flux of particle worldlines through the boundaries of each sub-domain, while the right hand side counts collisions within each sub-domain. As all boundaries are shared between sub-domains, the number of particle worldlines is conserved over the whole domain.
[…]
With these coordinate sets, the procedure for splitting the integration domain $Q$ into a continuous set of sub-domains can formally be achieved by taking the particle distribution function to be constant within each sub-domain of phase space. The total distribution is formed by the sum of these constant regions:
$\begin{split}f(\bm{x},\bm{p})=&\sum_{\alpha\beta\gamma\delta ijk}\frac{f_{\alpha\beta\gamma\delta ijk}}{p^{2}\Delta p_{i}\Delta u_{j}\Delta\phi_{k}}H_{\alpha}(t)H_{\beta}(x)\\ &\times H_{\gamma}(y)H_{\delta}(z)H_{i}(p)H_{j}(u)H_{k}(\phi),\end{split}$ (37)
where indices $\alpha,\beta,\gamma,\delta,i,j,$ and $k$ now refer to a discrete set of phase-space coordinates and $H$ is a boxcar function. As an example, for the $t$ coordinate the function $H$ is defined as
$\begin{split}H_{\alpha}(t)&=\Theta(t-t_{\alpha})-\Theta(t-t_{\alpha+1})\\ &=\begin{cases}1,&\text{for }t_{\alpha}<t<t_{\alpha+1},\\ h_{\alpha}^{-}&\text{for }t=t_{\alpha}\\ h_{\alpha}^{+}=1-h_{\alpha}^{-}&\text{for }t=t_{\alpha+1}\\ 0,&\text{for all other }t,\end{cases}\end{split}$ (38)
[…]
Through the action of Eq. (37), the continuous distribution function is replaced by a series of plateaux, each with a “height” of $f_{\alpha\beta\gamma\delta ijk}$ over a discrete sub-domain in phase space (see Fig. 6), and hence termed Distribution-In-Plateaux (DIP) [^9] It has been pointed out to the authors that this method of discretisation may be mapped to a zeroth-order discontinuous Galerkin method, which has been developed independently of this work, and further, has been shown to have good conservative properties for number density and energy (see, for example Cockburn and Shu, 2001; Gamba and Rjasanow, 2018; Gasparyan et al., 2022).. This nomenclature has similarities with that of the Particle-In-Cell (PIC), whereby the distribution function is considered as a sum over Dirac delta functions.
The factor of $p^{2}\Delta p_{i}\Delta u_{j}\Delta\phi_{k}$ in Eq. (37) is not strictly required, rather it has been introduced for convenience such that according to Eq. (28) the scalar number density $n$ of particles measured by the stationary observer, within a spatial sub-domain $X_{\beta\gamma\delta}$ and between times $t_{\alpha}$ and $t_{\alpha+1}$, is given by
$n=\sum_{ijk}f_{\alpha\beta\gamma\delta ijk},$ (39)
and as such $f_{\alpha\beta\gamma\delta ijk}$ can be recognised as the number density of particles within a momentum space sub-domain $P_{ijk}$.
The application of DIP form of the distribution function to the equations of phase-space transport and collision integrals simplifies their evaluation in terms of expressions that can be implemented computationally. However, this comes at the cost of these expressions becoming rather long. These long expressions can be found in full in Appendices A and C, which will be referred to in the following sections.
4.1 DIP Applied to Collision Integrals
Before describing how DIP simplifies the larger problem of phase-space transport, it is useful to consider how it simplifies the problem of evaluating collision integrals. For a binary collision $\mathfrak{ab\rightarrow cd}$, the gain of particles of type-$\mathfrak{c}$ is given by the collision integral (Eq. (36)):
$\begin{split}\int\bm{C}(\bm{x},\bm{p}_{\mathfrak{c}})=&\int f_{\mathfrak{a}}(\bm{x},\bm{p}_{\mathfrak{a}})f_{\mathfrak{b}}(\bm{x},\bm{p}_{\mathfrak{b}})\\ &\times G_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}\mathrm{d}^{3}p_{\mathfrak{a}}\mathrm{d}^{3}p_{\mathfrak{b}}\mathrm{d}^{3}p_{\mathfrak{c}}\bm{\chi},\end{split}$ (40)
where $G_{12\rightarrow 34}$ is the “gain term” given by Eq. (51) in Appendix A. The integration domain for this term is the phase space of particle $\mathfrak{c}$ and the momentum space of particles $\mathfrak{a}$ and $\mathfrak{b}$. By deploying DIP, the value of the distribution functions of particles $\mathfrak{a},\mathfrak{b}$, and $\mathfrak{c}$ become constant within a sub-domains of phase space and, therefore, can be taken outside integrals over those sub-domains. Consider a momentum-space sub-domain for particle $\mathfrak{c}$ to be $P_{ijk}=[p_{i},p_{i+1}]\times[u_{j},u_{j+1}]\times[\phi_{k},\phi_{k+1}]$, with DIP deployed, the collision integral (Eq. (40)) over this sub-domain becomes:
$\begin{split}\int^{t_{\alpha+1}}_{t_{\alpha}}\int_{X_{\beta\gamma\delta}}&\int_{P_{ijk}}\bm{C}(\bm{x},\bm{p}_{\mathfrak{c}})=\sum_{lmnopq}f_{\mathfrak{a},\alpha\beta\gamma\delta lmn}\\ &\times f_{\mathfrak{b},\alpha\beta\gamma\delta opq}G_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d},ijklmnopq}\mathcal{V}_{\alpha\beta\gamma\delta},\end{split}$ (41)
[…]
The application of DIP to collision integrals presents some immediate advantages of using DIP over PIC. As all collision array terms may be pre-evaluated and reused over multiple simulations of phase-space transport, there is a reduction in computational intensity compared to the calculation of collisions during transport—especially when the number of particles (collisions) is large. In addition, it provides uniform sampling of collisions over the entire phase-space domain—independent of the number of particles in the system.
4.2 DIP Applied to Transport
Consider an arbitrary but stationary spacetime, with a metric tensor given by:
$\begin{split}\bm{g}=&\left(-A^{2}+\frac{E^{2}}{B^{2}}+\frac{F^{2}}{C^{2}}+\frac{G^{2}}{D^{2}}\right)\bm{\mathrm{d}}t\otimes\bm{\mathrm{d}}t\\ &+B^{2}\bm{\mathrm{d}}x\otimes\bm{\mathrm{d}}x+C^{2}\bm{\mathrm{d}}y\otimes\bm{\mathrm{d}}y+D^{2}\bm{\mathrm{d}}z\otimes\bm{\mathrm{d}}z\\ &+2E\bm{\mathrm{d}}t\otimes\bm{\mathrm{d}}x+2F\bm{\mathrm{d}}t\otimes\bm{\mathrm{d}}y+2G\bm{\mathrm{d}}t\otimes\bm{\mathrm{d}}z,\end{split}$ (42)
Being stationary, the functions $A,B,C,D,E,F,$ and $G$ are functions of a subset of the spatial coordinates $x^{\iota}$ only. The functions $E,F,G$ form what is conventionally known as the shift vector $\beta_{\iota}=(E,F,G)$ with the function $A$ being the lapse function. The metric given by Eq. (42) could refer to a wide range of metrics commonly applied to astrophysical settings, an example subset of which are given in Appendix B.
Figure 7: Simplified schematic of how DIP applies to phase-space transport (here phase space has been reduced to two dimensions). The action of Distribution-In-Plateaux Eq. (37), is to discretise phase space into sub-domains in which the distribution function $f$ takes a single value ($f_{ij}$ here). The left hand side of the transport equation Eq. (10) $\int_{\partial Q}f\bm{\omega}$ is then an integral over the boundaries of these sub-domains, which generates the fluxes between sub-domains ($\mathcal{A}$ and $\mathcal{B}$ here). For this two dimensional example, the transport equation for $f_{22}$ takes the form $\int_{\partial Q_{22}}f\bm{\omega}=\left(h^{-}_{32}f_{32}+h^{+}_{22}f_{22}\right)\mathcal{A}^{+}+\left(h^{-}_{22}f_{22}+h^{+}_{12}f_{12}\right)\mathcal{A}^{-}+\left(h^{-}_{23}f_{23}+h^{+}_{22}f_{22}\right)\mathcal{B}^{+}+\left(h^{-}_{22}f_{22}+h^{+}_{21}f_{21}\right)\mathcal{B}^{-}$, where the $h$ terms (Eq. (38)) dictate the value of the distribution function on sub-domain boundaries.
The spacetime volume element (Eq. (42)) of the metric given by Eq. (42) has the simple form:
$\chi_{0123}=ABCD,$ (43)
and a natural coordinate transform to the tetrad of a stationary (Eulerian) observer $\bm{n}=-A\bm{\mathrm{d}}t$, could be written:
[…]
where $X^{\alpha}=(0,\frac{X_{1}}{B},\frac{X_{2}}{C},\frac{X_{3}}{D})$, $Y^{\alpha}=(0,\frac{Y_{1}}{B},\frac{Y_{2}}{C},\frac{Y_{3}}{D})$, and $X^{\alpha}=(0,\frac{Z_{1}}{B},\frac{Z_{2}}{C},\frac{Z_{3}}{D})$ are a set of mutually orthonormal vectors, which are also orthogonal to $\bm{n}$. It is often standard to align $\bm{X},\bm{Y},\text{ and }\bm{Z}$ to the $\{x,y,z\}$ coordinate directions, however, they have here been left general as it can be useful to align them to some other set of vectors e.g. the local magnetic and electric field directions (Chael et al., 2023; Gelles et al., 2025; Tsunetoe et al., 2025).
With $\bm{n}$ and $e_{a}^{~\alpha}$ defined, the integral form of the transport equation Eq. (10) can be evaluated using DIP over a phase space sub-domain $X_{\alpha\beta\gamma\delta}\times P_{ijk}=[t_{\alpha},t_{\alpha+1}]\times[x_{\beta},x_{\beta+1}]\times[y_{\gamma},y_{\gamma+1}]\times[z_{\delta},z_{\delta+1}]\times[p_{i},p_{i+1}]\times[u_{j},u_{j+1}]\times[\phi_{k},\phi_{k+1}]$ to give Eqs. (65) to (75) (found in Appendix C), whereby terms on the left hand side of Eq. (10) are converted into fluxes at the boundaries of the sub-domain (see Fig. 7) and those on the right hand side have been converted into a series of collision arrays. These equations form a numerical evolution scheme for all particle distribution functions, including external forces and collisions between particles. Such a scheme is conservative in terms of the total number of particles in the system (in the absence of emissive processes which are free to add or subtract particles from the system), as fluxes are always balanced between sub-domains, i.e. any particles leaving one sub-domain of phase space enter another.
As an example of how these fluxes are generated, consider the specific case of Schwarzchild geometry with coordinates $x^{\alpha}=\{t,r,\theta,\psi\}$ (see Appendix B, Eq. (62)). In this geometry $\bm{n}=-\sqrt{1-\frac{r_{s}}{r}}\bm{\mathrm{d}}t$, where $r_{s}$ is the Schwarzchild radius, and the coordinate transformation Eq. (44) may be given by
$e_{a}^{~\alpha}=\begin{pmatrix}\frac{1}{\sqrt{1-\frac{r_{s}}{r}}}&0&0&0\\ 0&\sqrt{1-\frac{r_{s}}{r}}&0&0\\ 0&0&\frac{1}{r}&0\\ 0&0&0&\frac{1}{r\sin\theta}\end{pmatrix}$, (45)
assuming the orthonormal basis vectors $\bm{X},\bm{Y}$, and $\bm{Z}$ are aligned to the global coordinate directions $\bm{e}_{r},\bm{e}_{\theta}$, and $\bm{e}_{\phi}$.
Using these definitions for the Eulerian observer $\bm{n}$ and coordinate transformations $e_{a}^{~\alpha}$, the fluxes through phase space sub-domain boundaries and the spacetime volume element (Eqs. (65) to (75)) can be evaluated analytically. For example, the flux elements $\mathcal{B}^{+}_{\alpha\beta\gamma\delta ijk}$ (Eq. (68)) through boundaries of constant $r$ are given by Eq. (LABEL:eqn:_Bp_flux_Schwarzchild).
Eq. (LABEL:eqn:_Bp_flux_Schwarzchild) is not a particularly elegant expression but importantly it, and all other flux expressions, are independent of the distribution function and any state variables of the system. Therefore all terms relating to transport, fluxes and collisions, are pre-computable as fixed arrays.
{widetext}
$\begin{split}&\mathcal{B}^{+}_{\alpha\beta\gamma\delta ijk}=-\frac{1}{2}\left(t_{\alpha+1}-t_{\alpha}\right)\left(r^{3/2}_{\beta}\sqrt{r_{\beta}-r_{s}}\right)\left(\cos\theta_{\gamma+1}-\cos\theta_{\gamma}\right)\left(\psi_{\delta+1}-\psi_{\delta}\right)\left(\sqrt{m^{2}+p^{2}_{i+1}}-\sqrt{m^{2}+p^{2}_{i}}\right)\\ &\times\left(u_{j+1}\sqrt{1-u_{j+1}^{2}}-2\text{arccot}\left(\frac{u_{j+1}-1}{\sqrt{1-u_{j+1}^{2}}}\right)-u_{j}\sqrt{1-u_{j}^{2}}+2\text{arccot}\left(\frac{u_{j}-1}{\sqrt{1-u_{j}^{2}}}\right)\right)\left(\sin\phi_{k+1}-\sin\phi_{k}\right).\end{split}$ (46)
5 Conclusion
This is the first paper in a series describing a novel mesoscopic framework for the general transport of particle distributions, with the aim of application to astrophysical systems.
In this paper, the transport of the particle distribution through phase space has been described in a general manner. Continuous forces and discrete particle interactions have been included (Sections 2 and 3).
This general description is independent of the underlying spacetime and agnostic to the types of particle interactions involved.
A novel approach of phase space discretisation of the particle distribution function, Distribution-In-Plateaux (Section 4), has been described that allows for the numerical evaluation of this transport in a conservative form.
[…]
It is planned to apply the DIPLODOCUS framework to AGN jet modelling in the future. In particular, the added dimensionality allowed by the relaxations of various assumptions on the particle distributions and jet geometries will be exploited. This will allow their effects on emission spectra to be examined and may provide insight into the jet content and particle acceleration mechanisms.
Acknowledgements
C. N. Everett acknowledges a Science and Technologies Facilities Council (STFC) studentship ST/X508664/1. G. Cotter acknowledges support from STFC grants ST/V006355/1, ST/V001477/1, and ST/S002952/1, and from Exeter College, Oxford. The authors would also like to thank James Matthews, Marc Klinger-Plaisier, Sera Markoff (and their research group), Pedro Ferreira, and Dion Everett for their helpful insights during various phases of this work.
Data Availability
No data were created or analysed in this article.
Appendix A Collision Integral for Binary and Emissive interactions
A.1 Binary Interactions
Consider the reversible binary (scattering) interaction $\mathfrak{a}\mathfrak{b}\rightleftharpoons\mathfrak{c}\mathfrak{d}$. The collision integral for a particle of type-$\mathfrak{c}$ in this interaction is given by Eq. (36):
$\bm{C}(\bm{x},\bm{p}_{\mathfrak{c}})=\bm{N}_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}-\bm{N}_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d}}.$ (47)
$\bm{N}_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}$ is the number of worldlines created by the forward interaction $\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}$ in this reversible process, hence giving the number of particle type-$\mathfrak{c}$ (and type-$\mathfrak{d}$) worldlines started by that interaction within a volume of phase space. Using Eq. (35) this may be written
$\bm{N}_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}=\delta^{(4)}(\bm{p}_{\mathfrak{a}}+\bm{p}_{\mathfrak{b}}-\bm{p}_{\mathfrak{c}}-\bm{p}_{\mathfrak{d}})\left|M_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}\right|^{2}f_{\mathfrak{a}}(\bm{x},\bm{p}_{\mathfrak{a}})f_{\mathfrak{b}}(\bm{x},\bm{p}_{\mathfrak{b}})\frac{\bm{\pi}_{\mathfrak{a}}\bm{\pi}_{\mathfrak{b}}}{(1+\delta_{\mathfrak{a}\mathfrak{b}})}\frac{\bm{\pi}_{\mathfrak{c}}\bm{\pi}_{\mathfrak{d}}}{(1+\delta_{\mathfrak{c}\mathfrak{d}})}\bm{\chi}.$ (48)
Three of the Dirac delta functions in above equation can be removed by integrating over the momentum states of the other output particle in the interaction, i.e. states of type-$\mathfrak{d}$ particles
$\bm{N}_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}=\delta(p^{0}_{\mathfrak{a}}+p^{0}_{\mathfrak{b}}-p^{0}_{\mathfrak{c}}-p^{0}_{\mathfrak{d}})\left|M_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}\right|^{2}f_{\mathfrak{a}}(\bm{x},\bm{p}_{\mathfrak{a}})f_{\mathfrak{b}}(\bm{x},\bm{p}_{\mathfrak{b}})\frac{\mathrm{d}^{3}p_{\mathfrak{a}}\mathrm{d}^{3}p_{\mathfrak{b}}}{p^{0}_{\mathfrak{a}}p^{0}_{\mathfrak{b}}(1+\delta_{\mathfrak{a}\mathfrak{b}})}\frac{\mathrm{d}^{3}p_{\mathfrak{c}}}{p^{0}_{\mathfrak{c}}p^{0}_{\mathfrak{d}}(1+\delta_{\mathfrak{c}\mathfrak{d}})}\bm{\chi},$ (49)
where $p^{0}_{\mathfrak{d}}$ is now a function of the other particle momenta given by energy conservation $p^{0}_{\mathfrak{d}}=p^{0}_{\mathfrak{a}}+p^{0}_{\mathfrak{b}}-p^{0}_{\mathfrak{c}}$ and note the shorthand for the momentum space volume element $\mathrm{d}^{3}p/p^{0}=(p)^{2}\mathrm{d}p\mathrm{d}u\mathrm{d}\phi/p^{0}=\bm{\mathrm{d}}p^{1}\wedge\bm{\mathrm{d}}p^{2}\wedge\bm{\mathrm{d}}p^{3}/p^{0}$.
For binary collisions, the rate is typically measured using a differential cross section. For generality of reference frame, it is best to use the Lorentz invariant differential cross section $\frac{\mathrm{d}\sigma_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}}{\mathrm{d}T}$, which is a function of the standard Mandelstam variables $S,T\text{ and }U$. This differential cross section can be related to the scattering matrix (Eq. (33)) by
$\left|M_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}\right|^{2}=\frac{\mathcal{F}_{\mathfrak{a}\mathfrak{b}}^{2}}{\pi}\frac{\mathrm{d}\sigma_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}}{\mathrm{d}T},$ (50)
where $\mathcal{F}_{\mathfrak{a}\mathfrak{b}}=\frac{1}{2}\sqrt{\left(S-(m_{\mathfrak{a}}+m_{\mathfrak{b}})^{2}\right)\left(S-(m_{\mathfrak{a}}-m_{\mathfrak{b}})^{2}\right)}$ is a Lorentz scalar related to the incident flux of particles of type-$\mathfrak{a}$ and type-$\mathfrak{b}$.
The internal terms of Eq. (49) may then be collected into the expression
$\bm{N}_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}=f_{\mathfrak{a}}(\bm{x},\bm{p}_{\mathfrak{a}})f_{\mathfrak{b}}(\bm{x},\bm{p}_{\mathfrak{b}})G_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}\mathrm{d}^{3}p_{\mathfrak{a}}\mathrm{d}^{3}p_{\mathfrak{b}}\mathrm{d}^{3}p_{\mathfrak{c}}\bm{\chi},\quad G_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}=\frac{\delta(p^{0}_{\mathfrak{a}}+p^{0}_{\mathfrak{b}}-p^{0}_{\mathfrak{c}}-p^{0}_{\mathfrak{d}})}{(1+\delta_{\mathfrak{a}\mathfrak{b}})(1+\delta_{\mathfrak{c}\mathfrak{d}})}\frac{\mathcal{F}_{\mathfrak{a}\mathfrak{b}}^{2}}{\pi p^{0}_{\mathfrak{a}}p^{0}_{\mathfrak{b}}p^{0}_{\mathfrak{c}}p^{0}_{\mathfrak{d}}}\frac{\mathrm{d}\sigma_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}}{\mathrm{d}T},$ (51)
with $G_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}$ being denoted the “gain term” of particles of type-$\mathfrak{c}$ (and type-$\mathfrak{d}$) for the forward reaction.
For the reverse process $\bm{N}_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d}}$, Eq. (53) can be altered and simplified by acknowledging that now both output states (now particles of type-$\mathfrak{a}$ and type-$\mathfrak{b}$) can be integrated over, leading to
$\bm{N}_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d}}=f_{\mathfrak{c}}(\bm{x},\bm{p}_{\mathfrak{c}})f_{\mathfrak{d}}(\bm{x},\bm{p}_{\mathfrak{d}})L_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d}}\mathrm{d}^{3}p_{\mathfrak{c}}\mathrm{d}^{3}p_{\mathfrak{d}}\bm{\chi},\quad L_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d}}=\frac{1}{\left(1+\delta_{\mathfrak{c}\mathfrak{d}}\right)}\frac{\mathcal{F}_{\mathfrak{c}\mathfrak{d}}\sigma_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d}}}{p^{0}_{\mathfrak{c}}p^{0}_{\mathfrak{d}}},$ (52)
where $\sigma_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d}}$ is the Lorentz invariant total cross section for the reverse process $\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d}$, which typically involves the factor of $1/(1+\delta_{\mathfrak{a}\mathfrak{b}})$ in its evaluation, and $L_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d}}$ is the “loss term” for particles of type-$\mathfrak{c}$ (and type-$\mathfrak{d}$) for the reverse reaction.
A.2 Binary Interactions in DIP Form
Under the assumption that the distribution function takes the form granted by DIP, Eq. (37), the gain and loss terms Eqs. (51) and (52) may be integrated over sub-domains of the momentum space of the particles involved to give the “gain array” components:
$G_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d},ijklmnopq}=\int_{P_{ijk}}\int_{P_{lmn}}\int_{P_{opq}}G_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d}}\frac{\mathrm{d}^{3}p_{\mathfrak{b}}}{(p_{\mathfrak{b}})^{2}\Delta p_{\mathfrak{b},o}\Delta u_{\mathfrak{b},p}\Delta\phi_{\mathfrak{b},q}}\frac{\mathrm{d}^{3}p_{\mathfrak{a}}}{(p_{\mathfrak{a}})^{2}\Delta p_{\mathfrak{a},l}\Delta u_{\mathfrak{a},m}\Delta\phi_{\mathfrak{a},n}}\mathrm{d}^{3}p_{\mathfrak{c}}$ (53)
and “loss array” components
$L_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d},ijklmn}=\int_{P_{ijk}}\int_{P_{lmn}}L_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d}}\frac{\mathrm{d}^{3}p_{\mathfrak{d}}}{(p_{\mathfrak{d}})^{2}\Delta p_{\mathfrak{d},l}\Delta u_{\mathfrak{d},m}\Delta\phi_{\mathfrak{d},n}}\frac{\mathrm{d}^{3}p_{\mathfrak{c}}}{(p_{\mathfrak{c}})^{2}\Delta p_{\mathfrak{c},i}\Delta u_{\mathfrak{c},j}\Delta\phi_{\mathfrak{c},k}}$ (54)
[…]
A.3 Emissive Interactions
Consider an emissive interaction where one particle “decays” into two $\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}$. The collision integral for a particle of type-$\mathfrak{c}$ in this interaction is given by Eq. (36):
$\bm{C}(\bm{x},\bm{p}_{\mathfrak{c}})=\bm{N}_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}},$ (55)
$\bm{N}_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}}$ is the number of worldline of particles of type-$\mathfrak{c}$ created by the emissive interaction for $\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}$. Using Eq. (35) this may be written:
$\bm{N}_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}}=\delta^{(4)}(\bm{p}_{\mathfrak{a}}-\bm{p}_{\mathfrak{b}}-\bm{p}_{\mathfrak{c}})\left|M_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}}\right|^{2}f_{\mathfrak{a}}(\bm{x},\bm{p}_{\mathfrak{a}})\bm{\pi}_{\mathfrak{a}}\frac{\bm{\pi}_{\mathfrak{b}}\bm{\pi}_{\mathfrak{c}}}{(1+\delta_{\mathfrak{b}\mathfrak{c}})}\bm{\chi}.$ (56)
Three of the Dirac delta functions in the above equation can be removed by integrating over the momentum states of the outgoing particle of type-$\mathfrak{b}$:
$\bm{N}_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}}=\frac{\delta(p^{0}_{\mathfrak{a}}-p^{0}_{\mathfrak{b}}-p^{0}_{\mathfrak{c}})\left|M_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}}\right|^{2}}{p^{0}_{\mathfrak{a}}p^{0}_{\mathfrak{b}}p^{0}_{\mathfrak{c}}(1+\delta_{\mathfrak{b}\mathfrak{c}})}f_{\mathfrak{a}}(\bm{x},\bm{p}_{\mathfrak{a}})\mathrm{d}^{3}p_{\mathfrak{a}}\mathrm{d}^{3}p_{\mathfrak{c}}\bm{\chi}=\frac{\mathrm{d}N_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}}}{\mathrm{d}x^{0}\mathrm{d}^{3}p_{\mathfrak{c}}}f_{\mathfrak{a}}(\bm{x},\bm{p}_{\mathfrak{a}})\mathrm{d}^{3}p_{\mathfrak{a}}\mathrm{d}^{3}p_{\mathfrak{c}}\bm{\chi},$ (57)
where $\frac{\mathrm{d}N_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}}}{\mathrm{d}x^{0}\mathrm{d}^{3}p_{\mathfrak{c}}}$ is the single particle emissivity of the interaction. This single particle emissivity is typically dependent on some external field such as the presence of a magnetic field in the case of synchrotron emissions.
A.4 Emissive Interactions in DIP Form
Under the assumptions that the distribution of the particles takes the form granted by DIP, Eq. (37), the gain array for an emissive interaction $\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}$ can be generated by integrating Eq. (57) over discrete sub-domains of momentum space:
$G_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c},ijklmn}=\int_{P_{ijk}}\int_{P_{lmn}}G_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}}\frac{\mathrm{d}^{3}p_{\mathfrak{a}}}{(p_{\mathfrak{a}})^{2}\Delta p_{\mathfrak{a},l}\Delta u_{\mathfrak{a},m}\Delta\phi_{\mathfrak{a},n}}\frac{\mathrm{d}^{3}p_{\mathfrak{c}}}{(p_{\mathfrak{c}})^{2}},\quad G_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}}=(p_{\mathfrak{c}})^{2}\frac{\mathrm{d}N_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c}}}{\mathrm{d}x^{0}\mathrm{d}^{3}p_{\mathfrak{c}}}$ (58)
Appendix B Metrics
A metric of the form Eq. (42), can express most standard astrophycially relevant metrics. A subset are presented below:
Cartesian Minkowski:
$x^{\alpha}=\{t,x,y,z\},~A^{2}=1,~B^{2}=1,~C^{2}=1,~D^{2}=1,~E=F=G=0.$ (59)
Spherical Minkowski:
$x^{\alpha}=\{t,r,\theta,\psi\},~A^{2}=1,~B^{2}=1,~C^{2}=r^{2},~D^{2}=r^{2}\sin^{2}\theta,~E=F=G=0.$ (60)
Cylindrical Minkowski:
$x^{\alpha}=\{t,\rho,\vartheta,z\},~A^{2}=1,~B^{2}=1,~C^{2}=\rho^{2},~D^{2}=1,~E=F=G=0.$ (61)
Schwarzschild:
$\quad x^{\alpha}=\{t,r,\theta,\psi\},~A^{2}=\left(1-\frac{r_{s}}{r}\right),~B^{2}=\left(1-\frac{r_{s}}{r}\right)^{-1},~C^{2}=r^{2},~D^{2}=r^{2}\sin^{2}\theta,~E=F=G=0.$ (62)
Kerr (Boyer-Lindquist Coordinates):
$\begin{split}x^{\alpha}&=\{t,r,\theta,\psi\},~A^{2}=\left(1-\frac{r_{s}}{r}\right),~B^{2}=\frac{\Sigma}{\Delta},~C^{2}=\Sigma,~D^{2}=\left(r^{2}+a^{2}+\frac{r_{s}ra^{2}}{\Sigma}\sin^{2}\theta\right)\sin^{2}\theta,~E=F=0,\\ &\quad\quad G=\frac{-r_{s}ra\sin^{2}\theta}{\Sigma},~\Sigma=r^{2}+a^{2}\cos^{2}\theta,~\Delta=r^{2}-rr_{s}+a^{2}.\end{split}$ (63)
Appendix C DIP Transport Equations
Under the assumption that the distribution function $f(\bm{x},\bm{p})$ for all particles takes the DIP form, Eq. (37), the integral form of the transport equation Eq. (10), for a particle of type-$\mathfrak{c}$ can be evaluated over a phase-space sub-domain $X_{\alpha\beta\gamma\delta}\times P_{ijk}=[t_{\alpha},t_{\alpha+1}]\times[x_{\beta},x_{\beta+1}]\times[y_{\gamma},y_{\gamma+1}]\times[z_{\delta},z_{\delta+1}]\times[p_{i},p_{i+1}]\times[u_{j},u_{j+1}]\times[\phi_{k},\phi_{k+1}]$ to give the discrete DIP transport equation:
$\begin{split}&\left(h^{-}_{\alpha+1}f_{\mathfrak{c},(\alpha+1)\beta\gamma\delta ijk}+h^{+}_{\alpha}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}\right)\mathcal{A}^{+}_{\alpha\beta\gamma\delta ijk}+\left(h^{-}_{\alpha}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}+h^{+}_{\alpha-1}f_{\mathfrak{c},(\alpha-1)\beta\gamma\delta ijk}\right)\mathcal{A}^{-}_{\alpha\beta\gamma\delta ijk}\\ &\quad\quad+\left(h^{-}_{\beta+1}f_{\mathfrak{c},\alpha(\beta+1)\gamma\delta ijk}-h^{+}_{\beta}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}\right)\mathcal{B}^{+}_{\alpha\beta\gamma\delta ijk}+\left(h^{-}_{\beta}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}-h^{+}_{\beta-1}f_{\mathfrak{c},\alpha(\beta-1)\gamma\delta ijk}\right)\mathcal{B}^{-}_{\alpha\beta\gamma\delta ijk}\\ &\quad\quad+\left(h^{-}_{\gamma+1}f_{\mathfrak{c},\alpha\beta(\gamma+1)\delta ijk}+h^{+}_{\gamma}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}\right)\mathcal{C}^{+}_{\alpha\beta\gamma\delta ijk}+\left(h^{-}_{\gamma}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}+h^{+}_{\gamma-1}f_{\mathfrak{c},\alpha\beta(\gamma-1)\delta ijk}\right)\mathcal{C}^{-}_{\alpha\beta\gamma\delta ijk}\\ &\quad\quad+\left(h^{-}_{\delta+1}f_{\mathfrak{c},\alpha\beta\gamma(\delta+1)ijk}+h^{+}_{\delta}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}\right)\mathcal{D}^{+}_{\alpha\beta\gamma\delta ijk}+\left(h^{-}_{\delta}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}+h^{+}_{\delta-1}f_{\mathfrak{c},\alpha\beta\gamma(\delta-1)ijk}\right)\mathcal{D}^{-}_{\alpha\beta\gamma\delta ijk}\\ &\quad\quad+\left(\frac{h^{-}_{i+1}f_{\mathfrak{c},\alpha\beta\gamma\delta(i+1)jk}}{\Delta p_{\mathfrak{c},i+1}}+\frac{h^{+}_{i}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}}{\Delta p_{\mathfrak{c},i}}\right)\mathcal{I}^{+}_{\alpha\beta\gamma\delta ijk}+\left(\frac{h^{-}_{i}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}}{\Delta p_{\mathfrak{c},i}}+\frac{h^{+}_{i-1}f_{\mathfrak{c},\alpha\beta\gamma\delta(i-1)jk}}{\Delta p_{\mathfrak{c},i-1}}\right)\mathcal{I}^{-}_{\alpha\beta\gamma\delta ijk}\\ &\quad\quad+\left(\frac{h^{-}_{j+1}f_{\mathfrak{c},\alpha\beta\gamma\delta i(j+1)k}}{\Delta u_{\mathfrak{c},j+1}}+\frac{h^{+}_{j}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}}{\Delta u_{\mathfrak{c},j}}\right)\mathcal{J}^{+}_{\alpha\beta\gamma\delta ijk}+\left(\frac{h^{-}_{j}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}}{\Delta u_{\mathfrak{c},j}}+\frac{h^{+}_{j-1}f_{\mathfrak{c},\alpha\beta\gamma\delta i(j-1)k}}{\Delta u_{\mathfrak{c},j-1}}\right)\mathcal{J}^{-}_{\alpha\beta\gamma\delta ijk}\\ &\quad\quad+\left(\frac{h^{-}_{k+1}f_{\mathfrak{c},\alpha\beta\gamma\delta ij(k+1)}}{\Delta\phi_{\mathfrak{c},k+1}}+\frac{h^{+}_{k}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}}{\Delta\phi_{\mathfrak{c},k}}\right)\mathcal{K}^{+}_{\alpha\beta\gamma\delta ijk}+\left(\frac{h^{-}_{k}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}}{\Delta\phi_{\mathfrak{c},k}}+\frac{h^{+}_{k-1}f_{\mathfrak{c},\alpha\beta\gamma\delta ij(k-1)}}{\Delta\phi_{\mathfrak{c},k-1}}\right)\mathcal{K}^{-}_{\alpha\beta\gamma\delta ijk}\\ &=\sum_{\mathfrak{a}\mathfrak{b}\mathfrak{d}\in\text{particles}}\left(\sum_{lmnopq}G_{\mathfrak{a}\mathfrak{b}\rightarrow\mathfrak{c}\mathfrak{d},ijklmnopq}f_{\mathfrak{a},\alpha\beta\gamma\delta lmn}f_{\mathfrak{b},\alpha\beta\gamma\delta opq}-\sum_{lmn}L_{\mathfrak{a}\mathfrak{b}\leftarrow\mathfrak{c}\mathfrak{d},ijklmn}f_{\mathfrak{c},\alpha\beta\gamma\delta ijk}f_{\mathfrak{d},\alpha\beta\gamma\delta lmn}\right)\mathcal{V}_{\alpha\beta\gamma\delta}\\ &\quad\quad+\sum_{\mathfrak{a}\mathfrak{b}\in\text{particles}}\sum_{lmn}G_{\mathfrak{a}\rightarrow\mathfrak{b}\mathfrak{c},ijklmn}f_{\mathfrak{a},\alpha\beta\gamma\delta lmn}\mathcal{V}_{\alpha\beta\gamma\delta}\end{split}$ (64)
where the spacetime volume element is
$\mathcal{V}_{\alpha\beta\gamma\delta}=\int_{z_{\delta}}^{z_{\delta+1}}\int_{y_{\gamma}}^{y_{\gamma+1}}\int_{x_{\beta}}^{x_{\beta+1}}\int_{t_{\alpha}}^{t_{\alpha+1}}\chi_{0123}\mathrm{d}t\mathrm{d}x\mathrm{d}y\mathrm{d}z,$ (65)
the fluxes through hypersurfaces of constant time coordinate are:
$\mathcal{A}^{+}_{\alpha\beta\gamma\delta ijk}$ $=\int_{\phi_{k}}^{\phi_{k+1}}\int_{u_{j}}^{u_{j+1}}\int_{p_{i}}^{p_{i+1}}\int_{z_{\delta}}^{z_{\delta+1}}\int_{y_{\gamma}}^{y_{\gamma+1}}\int_{x_{\beta}}^{x_{\beta+1}}\left[p^{a}e_{a}^{~0}\frac{\chi_{0123}}{p^{0}}\right]_{t_{\alpha+1}}\frac{\mathrm{d}x\mathrm{d}y\mathrm{d}z\mathrm{d}p\mathrm{d}u\mathrm{d}\phi}{\Delta p_{i}\Delta u_{j}\Delta\phi_{k}},$ (66)
$\mathcal{A}^{-}_{\alpha\beta\gamma\delta ijk}$ $=-\mathcal{A}^{+}_{(\alpha-1)\beta\gamma\delta ijk}$, (67)
the fluxes through hypersurfaces of constant spatial coordinates are:
$\mathcal{B}^{+}_{\alpha\beta\gamma\delta ijk}$ $=\int_{\phi_{k}}^{\phi_{k+1}}\int_{u_{j}}^{u_{j+1}}\int_{p_{i}}^{p_{i+1}}\int_{t_{\alpha}}^{t_{\alpha+1}}\int_{z_{\delta}}^{z_{\delta+1}}\int_{y_{\gamma}}^{y_{\gamma+1}}\left[p^{a}e_{a}^{~1}\frac{\chi_{1230}}{p^{0}}\right]_{x_{\beta+1}}\frac{\mathrm{d}y\mathrm{d}z\mathrm{d}t\mathrm{d}p\mathrm{d}u\mathrm{d}\phi}{\Delta p_{i}\Delta u_{j}\Delta\phi_{k}},\quad$ (68)
$\mathcal{C}^{+}_{\alpha\beta\gamma\delta ijk}$ $=\int_{\phi_{k}}^{\phi_{k+1}}\int_{u_{j}}^{u_{j+1}}\int_{p_{i}}^{p_{i+1}}\int_{x_{\beta}}^{x_{\beta+1}}\int_{t_{\alpha}}^{t_{\alpha+1}}\int_{z_{\delta}}^{z_{\delta+1}}\left[p^{a}e_{a}^{~2}\frac{\chi_{2301}}{p^{0}}\right]_{y_{\gamma+1}}\frac{\mathrm{d}z\mathrm{d}t\mathrm{d}x\mathrm{d}p\mathrm{d}u\mathrm{d}\phi}{\Delta p_{i}\Delta u_{j}\Delta\phi_{k}},\quad$ (69)
$\mathcal{D}^{+}_{\alpha\beta\gamma\delta ijk}$ $=\int_{\phi_{k}}^{\phi_{k+1}}\int_{u_{j}}^{u_{j+1}}\int_{p_{i}}^{p_{i+1}}\int_{y_{\gamma}}^{y_{\gamma+1}}\int_{x_{\beta}}^{x_{\beta+1}}\int_{t_{\alpha}}^{t_{\alpha+1}}\left[p^{a}e_{a}^{~3}\frac{\chi_{3012}}{p^{0}}\right]_{z_{\delta+1}}\frac{\mathrm{d}t\mathrm{d}x\mathrm{d}y\mathrm{d}p\mathrm{d}u\mathrm{d}\phi}{\Delta p_{i}\Delta u_{j}\Delta\phi_{k}},\quad$ (70)
$\mathcal{B}^{-}_{\alpha\beta\gamma\delta ijk}$ $=-\mathcal{B}^{+}_{\alpha(\beta-1)\gamma\delta ijk},\quad\mathcal{C}^{-}_{\alpha\beta\gamma\delta ijk}=-\mathcal{C}^{+}_{\alpha\beta(\gamma-1)\delta ijk},\quad\mathcal{D}^{-}_{\alpha\beta\gamma\delta ijk}=-\mathcal{D}^{+}_{\alpha\beta\gamma(\delta-1)ijk}$, (71)
the fluxes through constant momentum coordinates are:
$\mathcal{I}^{+}_{\alpha\beta\gamma\delta ijk}$ $=\int_{\phi_{k}}^{\phi_{k+1}}\int_{u_{j}}^{u_{j+1}}\int_{z_{\delta}}^{z_{\delta+1}}\int_{y_{\gamma}}^{y_{\gamma+1}}\int_{x_{\beta}}^{x_{\beta+1}}\int_{t_{\alpha}}^{t_{\alpha+1}}\left[\frac{\chi_{0123}}{p^{0}}R^{1}_{~a}\left(-\Gamma^{a}_{~bc}p^{b}p^{c}+mF^{a}\right)\right]_{p_{i+1}}\frac{\mathrm{d}t\mathrm{d}x\mathrm{d}y\mathrm{d}z\mathrm{d}u\mathrm{d}\phi}{\Delta u_{j}\Delta\phi_{k}},$ (72)
$\mathcal{J}^{+}_{\alpha\beta\gamma\delta ijk}$ $=\int_{p_{i}}^{p_{i+1}}\int_{\phi_{k}}^{\phi_{k+1}}\int_{z_{\delta}}^{z_{\delta+1}}\int_{y_{\gamma}}^{y_{\gamma+1}}\int_{x_{\beta}}^{x_{\beta+1}}\int_{t_{\alpha}}^{t_{\alpha+1}}\left[\frac{\chi_{0123}}{p^{0}}R^{2}_{~a}\left(-\Gamma^{a}_{~bc}p^{b}p^{c}+mF^{a}\right)\right]_{u_{j+1}}\frac{\mathrm{d}t\mathrm{d}x\mathrm{d}y\mathrm{d}z\mathrm{d}\phi\mathrm{d}p}{\Delta p_{i}\Delta\phi_{k}},$ (73)
$\mathcal{K}^{+}_{\alpha\beta\gamma\delta ijk}$ $=\int_{u_{j}}^{u_{j+1}}\int_{p_{i}}^{p_{i+1}}\int_{z_{\delta}}^{z_{\delta+1}}\int_{y_{\gamma}}^{y_{\gamma+1}}\int_{x_{\beta}}^{x_{\beta+1}}\int_{t_{\alpha}}^{t_{\alpha+1}}\left[\frac{\chi_{0123}}{p^{0}}R^{3}_{~a}\left(-\Gamma^{a}_{~bc}p^{b}p^{c}+mF^{a}\right)\right]_{\phi_{k+1}}\frac{\mathrm{d}t\mathrm{d}x\mathrm{d}y\mathrm{d}z\mathrm{d}p\mathrm{d}u}{\Delta p_{i}\Delta u_{j}},$ (74)
$\mathcal{I}^{-}_{\alpha\beta\gamma\delta ijk}$ $=-\mathcal{I}^{+}_{\alpha\beta\gamma\delta(i-1)jk},\quad\mathcal{J}^{-}_{\alpha\beta\gamma\delta ijk}=-\mathcal{J}^{+}_{\alpha\beta\gamma\delta i(j-1)k},\quad\mathcal{K}^{-}_{\alpha\beta\gamma\delta ijk}=-\mathcal{K}^{+}_{\alpha\beta\gamma\delta ij(k-1)}$, (75)
with $p^{a}=(p^{0},p\sqrt{1-u^{2}}\cos\phi,p\sqrt{1-u^{2}}\sin\phi,pu)$ in modified spherical coordinates, with the associated transformation for vector components
$R^{a}_{~b}=\begin{pmatrix}1&0&0&0\\ 0&\sqrt{1-u^{2}}\cos\phi&\sqrt{1-u^{2}}\sin\phi&u\\ 0&-\frac{u\sqrt{1-u^{2}}\cos\phi}{p}&-\frac{u\sqrt{1-u^{2}}\sin\phi}{p}&\frac{1-u^{2}}{p}\\ 0&-\frac{\sin\phi}{p\sqrt{1-u^{2}}}&\frac{\cos\phi}{p\sqrt{1-u^{2}}}&0\\ \end{pmatrix}$. (76)
The volume elements and fluxes (Eqs. (65) to (75)) are all integrals over phase-space sub-domains, which in most cases can be analytically integrated as the form of all the internal functions, connection coefficients, external forces etc. are known (see, for example, Eq. (LABEL:eqn:_Bp_flux_Schwarzchild)).